OpenStudy (dj3strella):

Question 3.3. Select the proper equation in point-slope form for the given point and slope. Point: Slope: (Points : 1)

OpenStudy (dj3strella):

y+4= -5/6(x+8) y+4= -5/6(x-8) y-4= -5/6(x+8) y-4= 5/6(x-8)

OpenStudy (anonymous):

y- y1 = m(x-x1) m = slope

OpenStudy (anonymous):

in this case i don't see the points nor the slope at you're question